Calculate the decrease in temperature when 2.00 L at 20.0 ◦C is com-
pressed to 1.00 L.
Answer #1

According to Charles' law,

The volume,V of an ideal gas at constant pressure is directly proportional to its absolute temperature,T.

Let V1 and T1 be the initial volume and temperature of a gas and V2 and T2 be its final volume and temperature respectively.

Then,

  

Here,

V1 = 2 L

V2 = 1L

T1 = 20 ℃ = 20 + 273 =293 K

T2 = (V1/V2) × T1

= (1/2) × 293

= 146.5 K

Decrease in temperature = T1 - T2

= 293 - 146.5

= 146.5 K